OpenStudy (anonymous):

Example A jeweler is using an alloy, a mixture of metals. He has 10 pounds of the alloy. 2% is the rare metal robium. He has another source which is 5% robium. He wishes to combine these mixtures to form a mixture that is 3% robium. Solution: 1. Identify the variables. (0.02)(10) The quantity of robium in material 1 (5%)(x) The quantity of robium in material 2 (0.03)(x+10) The quantity of the desired mixture 2. Write an equation. 5% x + (0.02)(10) = (0.03)(x+10) The equation required. 3. Perform the Solution. 0.05x + 0.2 = .03x + 0.3 = 0.02x = 0.1 4. Write the answer. x = 5 The desired amount.

OpenStudy (anonymous):

And i need help with these problem. An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How much (to the nearest tenth lb.) of the second alloy must be added to the first alloy to get a 12% mixture. ______ pounds

OpenStudy (anonymous):

(0.15)(20) (10%)(x) (0.12)(x+20)

OpenStudy (anonymous):

so far right???

OpenStudy (anonymous):

yess :)

OpenStudy (anonymous):

so next 10%x(0.15)(20)=(0.12)(x+20) right

OpenStudy (anonymous):

yess. seems like you have it figured out.

OpenStudy (anonymous):

im confuse on last part

OpenStudy (andriod09):

What do you need now Claire?

OpenStudy (anonymous):

perform the solution

OpenStudy (anonymous):

hrmmm i got .7 ?

OpenStudy (andriod09):

you mean solve the following: \[.1x(.15)(.20)=(.12)(x+20)?\]

OpenStudy (anonymous):

.10x+3=.12x+2.4=

OpenStudy (andriod09):

.10=.1 its 10, you don't need the 0, but it could be easier. to solve. \[.1x+3=.12x+2.4\]\[.1x+.6=.12x\]\[.6=.11x\]\[x=1.833\infty(3)\]

OpenStudy (anonymous):

so how did you get .6

OpenStudy (andriod09):

.06* my bad 1 sec. \[.1x+.06=.12x\][\.06=.11x\]\[1.833\infty(3)=x\] oddly,its the same answer.

OpenStudy (anonymous):

1.8 pounds???

OpenStudy (andriod09):

you could always use the percent bar method. that's easiest.

OpenStudy (anonymous):

so im confused how many pounds.

OpenStudy (andriod09):

A jeweler is using an alloy, a mixture of metals. He has 10 pounds of the alloy. 2% is the rare metal robium. He has another source which is 5% robium. He wishes to combine these mixtures to form a mixture that is 3% robium.